### v2

# Direction Estimation of Sound Sources Using Microphone Arrays and Signal Strength

Abstract
Sound-tracking refers to the process of determining the direction from which a sound originates, making it a fundamental component of sound source localization. This capability is essential in a variety of applications, including security systems, acoustic monitoring, and speaker tracking, where accurately identifying the direction of a sound source enables real-time responses, efficient resource allocation, and improved situational awareness. While sound-tracking is closely related to localization, it specifically focuses on identifying the direction of the sound source rather than estimating its exact position in space. Despite its utility, sound-tracking systems face several challenges, such as maintaining directional accuracy and precision, along with the need for sophisticated hardware configurations and complex signal processing algorithms. This paper presents a sound-tracking method using three electret microphones. We estimate the direction of a sound source using a lightweight method that analyzes signals from three strategically placed microphones. By comparing the average power of the received signals, the system infers the most probable direction of the sound. The results indicate that the power level from each microphone effectively determines the sound source direction. The proposed technique offers distinct advantages over existing sound-tracking methods. Our system employs a straightforward and cost-effective hardware design, ensuring simplicity and affordability in implementation. It achieves a localization error of less than 6 degrees and a precision of 98%. Additionally, its effortless integration with various systems makes it versatile and adaptable. Consequently, this technique presents a robust and reliable solution for sound-tracking and localization, with potential applications spanning diverse domains such as security systems, smart homes, and acoustic monitoring.
Index Terms:
Sound tracking, Sound source localization, Direction estimation, Microphone arrays
The source code of the project is available on our GitHub repository:
https://github.com/mahdi943/soundlocalization
I Introduction
The ability to detect sounds and determine their direction has become important in modern intelligent systems. This functionality supports various applications, such as smart home automation and interactive voice-based systems. For instance, identifying the direction of a sound can help locate a speaker in a room, trace the source of environmental noise, or respond appropriately to specific acoustic events [1]. As demand grows for systems that are more responsive to their surroundings, directional sound sensing has the potential as a key component in enhancing situational awareness and context-aware behavior.
Technology designed to locate and follow sound sources has been widely integrated into robotics [2], search-and-rescue operations [3], and surveillance [4], demonstrating its effectiveness in real-world scenarios [5]. Such advancements enable more intuitive interactions with devices, improving responsiveness to user commands. We believe directional audio detection is particularly beneficial in applications where pinpointing a sound’s origin enhances functionality. For instance, in smart homes, voice assistants like Amazon Alexa and Google Assistant utilize this capability to determine a user’s location within a room, allowing for more effective responses. Similarly, adaptive audio systems in home theaters use spatial awareness to enhance sound distribution, creating immersive listening experiences tailored to the user’s position.
While sound localization generally refers to estimating the exact position of a sound source in space, sound tracking specifically focuses on determining the direction from which the sound originates. Localization aims to provide precise spatial coordinates, which often requires complex sensor arrays and extensive computation. In contrast, sound tracking emphasizes directional information, enabling real-time responsiveness with fewer sensors and simpler processing. This distinction allows sound tracking to serve as a lightweight, cost-effective alternative in scenarios where directionality is sufficient for the task at hand.
Many existing sound localization methods rely on complex hardware setups, intensive signal processing, and high computational demands. While these approaches offer high accuracy, they are often impractical for lightweight, low-cost applications. To address this, we propose a simple and efficient method for detecting the direction of sound sources using a minimal sensor setup, designed to operate with limited processing resources.
This work focuses on enabling directional sound-tracking for use in responsive systems such as smart home devices, interactive robots, and compact automation solutions. By estimating the direction of incoming sounds in real time, the proposed method supports adaptive behavior in resource-constrained environments, offering a practical alternative to more computationally intensive localization techniques.
Our approach employs a vector-based technique to accurately detect sound direction by computing the sum of sound signal vectors arranged in a rectangular configuration. For instance, if a sound originates from the left, the vector representation will reflect this directional characteristic. This capability is crucial for swiftly identifying potential threats in security systems and enhancing the responsiveness of smart home devices by dynamically adjusting settings based on user positions.
While machine learning and beamforming techniques are prevalent in the field of sound localization, our method distinguishes itself through its simplicity and affordability. By utilizing readily available Arduino boards and electret microphones, we simplify the assembly and configuration process, thereby enhancing accessibility for various applications. This approach holds great promise for improving systems that rely on precise sound source detection and tracking, especially in environments where budget constraints or deployment speed are critical considerations.
II Related Work
II-A Sound Source Localization (SSL)
Sound source localization (SSL) is a core component in many applications, including source separation [6], automatic speech recognition (ASR) [7], speech enhancement [8], noise control [9], and room acoustics analysis [10]. SSL typically aims to determine the position or direction of a sound source using sensor data.
Classical approaches to SSL include Time Difference of Arrival (TDOA), which computes the delay in signal arrival between multiple microphones, and phase-based techniques that leverage phase shifts of the audio signal across sensors [11, 12]. Beamforming methods also play a central role in SSL, using microphone arrays to estimate direction based on signal alignment over time [13, 14].
Other localization strategies involve Bayesian filtering [15, 16], triangulation [17, 18], and direction-surface intersection models [19, 20], all offering increased precision at the cost of complexity.
Microphone arrays are frequently employed for their spatial resolution capabilities, although systems with large arrays or high sampling rates can become computationally intensive [21]. Recent research has explored machine learning approaches to classify sound direction using raw or pre-processed signal data [22, 23].
II-B Sound Source Tracking
In dynamic environments, tracking the direction of a moving sound source over time is essential. Traditional SSL methods have been extended to enable real-time tracking using Kalman filters, particle filters, or hybrid probabilistic models [24]. These models can update the sound source’s estimated direction continuously, improving robustness to noise and source movement.
Sound-based sensing also plays a significant role in human-centered applications, such as speaker tracking for smart environments [25], productivity tools [26, 27], and acoustic monitoring in mobile and wearable devices [22].
One relevant example is intensity-based tracking, where the direction is inferred by comparing sound intensity differences across microphones [28]. While less complex than TDOA or beamforming, these methods may offer faster processing at the expense of some spatial accuracy.
II-C Summary and Motivation
While traditional SSL and tracking methods offer high accuracy, they often require large microphone arrays, complex signal processing algorithms, and high computational power—making them less ideal for lightweight or low-cost systems.
In contrast, our work focuses on a simple, intensity-based method that relies on only three microphones and does not require time delay or phase computations. This enables a more practical, scalable, and cost-effective solution for indoor directional sound-tracking, suitable for applications such as smart homes, speaker-following systems, and acoustic event detection.
III Approach
In this section, we describe the hardware and software setup used to implement our sound localization system. We then provide a detailed explanation of the components, the underlying methodology, and the measurement process involved in evaluating the system’s performance.
III-A Methodology
In this section, we explain how our approach determines the direction of a sound source by analyzing the intensity of sound signals captured by three microphones. The microphones are placed at 0°, 120°, and 240°, measured counterclockwise from a defined 0° reference axis aligned with the x-axis of the device. When the sound intensity detected by any of the microphones exceeds a predefined threshold (Algorithm 1, line 3), the system begins collecting a fixed number of amplitude samples from all three microphones (line 6). These samples reflect the sound signal’s strength over time and are used in the subsequent direction estimation process.
1 Input: Sound signal $S$, number of samples NN, threshold $T$
2 Output: Sound angle, sound magnitude, servo motor position
3if $S>T$ then
4 foreach microphone $M_{i}$ do
5    // Collect and average samples
6    $Array_{i}\leftarrow$ Collect NN samples from $S$
7    $Avg_{i}\leftarrow$ Average of $Array_{i}$
8
9    // Convert average signals to polar form
10    $V_{i}\leftarrow$ Polar vector from $Avg_{i}$
11    $R_{i}\leftarrow$ Rectangular form of $V_{i}$
12
13   end foreach
14
15 // Calculate the resultant vector
16 $X\leftarrow\sum R_{i}[0]$// Sum of x-components
17 $Y\leftarrow\sum R_{i}[1]$// Sum of y-components
18
19 // Compute final magnitude and angle
20 $ResultMag\leftarrow\sqrt{X^{2}+Y^{2}}$
21 $ResultAngle\leftarrow\text{atan2}(Y,X)$
22
23 // Showing the position with servo motor
24 $Servo\leftarrow ServoPos$
25 end if
Algorithm 1 Sound Localization Algorithm
[…]
$\alpha=(120^{\circ},\text{PowerAvg}_{1})$ (1)
$\beta=(0^{\circ},\text{PowerAvg}_{2})$ (2)
$\gamma=(240^{\circ},\text{PowerAvg}_{3})$ (3)
These polar vectors are converted to rectangular (Cartesian) form using standard trigonometric conversion, resulting in vectors $R=(x,y)$, where $x=r\cdot\cos(\theta)$ and $y=r\cdot\sin(\theta)$.
$R_{1}=(r_{\alpha}\cdot\cos(\theta_{\alpha}),\ r_{\alpha}\cdot\sin(\theta_{\alpha}))$ (4)
$R_{2}=(r_{\beta}\cdot\cos(\theta_{\beta}),\ r_{\beta}\cdot\sin(\theta_{\beta}))$ (5)
$R_{3}=(r_{\gamma}\cdot\cos(\theta_{\gamma}),\ r_{\gamma}\cdot\sin(\theta_{\gamma}))$ (6)
The three rectangular vectors are summed component-wise $R_{\text{total}}=(X,Y)$ to compute the resulting direction vector, where:
$X=\sum_{i=1}^{3}R_{x_{i}}$ (7)
$Y=\sum_{i=1}^{3}R_{y_{i}}$ (8)
The final direction of the sound source is determined by converting the resultant rectangular vector back to polar coordinates. The magnitude and angle are computed as follows:
$\text{ResultMag}=\sqrt{X^{2}+Y^{2}}$ (9)
$\text{ResultAngle}=\text{atan2}(Y,X)$ (10)
The atan2 function returns the angle (in radians) between the positive $x$-axis and the point $(X,Y)$, accounting for the correct quadrant of the angle. This ensures an accurate directional result in the full $\pi,\pi]$ range.
As illustrated in Figure 1, the black vectors represent the average intensity from each microphone, and the red vector shows their summation, indicating the estimated direction of the sound source. This method allows real-time directional estimation using minimal hardware and straightforward computations.
Figure 1: Example showing the magnitudes of signals from three different microphones in black, and their vector sum in red, representing the resulting sound direction.
III-B Implementation
The implemented sound-tracking system consists of three electret microphones arranged on a circular surface at equal angular intervals (0°, 120°, and 240°), each positioned 15 cm from the center. These microphones capture sound signals from different directions and are connected to LM358 operational amplifiers to boost the signal strength. The amplified outputs are fed into an Arduino board, which processes the signals and determines the direction of the incoming sound. A servo motor, also connected to the Arduino, rotates an indicator shaft toward the detected sound direction. Figure 2 illustrates the complete hardware setup, which is designed to be low-cost, efficient, and suitable for compact applications.
Figure 2: Circuit diagram represents the location of microphones and the servo motor in the center. Each microphone is 120 degrees from the next microphone, which shows on the green axis.
IV Results
IV-A Experimental Setup and Design
To evaluate the performance of our sound-tracking system, we conducted 60 experiments using a 2-second audio clip played from a cellphone. The phone was positioned 35 cm from the center of the robot—close enough to ensure that all microphones received the signal, yet far enough to avoid saturation or directional bias. The phone was placed at different angles (20° and 120°) on a circular platform, and the system estimated the direction of the sound using the algorithm described earlier.
Each sound event triggered the collection of a fixed number of samples from the three microphones. The servo motor’s directional tracking accuracy was then evaluated by converting the resulting vectors into rectangular coordinates using Equations 7 and 8. These were plotted as polar scatter plots (scatterpolar) to visually represent the direction and magnitude of the resulting vectors in (X,Y) format.
IV-B Vector Visualization and Interpretation
For each angle setting (120° and 20°), 30 tests were conducted. The vectors generated from each test were visualized in Figures 3 and 4, respectively. Each dot in the scatterplot represents one test instance. The concentration of points along distinct directions indicates consistent directional estimation. Variations in the distance of points from the origin reflect differences in signal strength—stronger signals appearing farther from the center, and weaker ones appearing closer.
Figure 3: Detecting sound source angles and magnitude at 120 degrees
Figure 4: Detecting sound source angles and magnitude at 0 degrees
Variability in signal magnitude arises due to environmental noise and the inherent sensitivity differences among microphones. These factors contribute to minor deviations from ideal alignment in the plotted vectors but do not significantly affect the overall directional accuracy.
IV-C Accuracy and Precision Analysis
To ensure robustness, a 6% trimming was applied to the sorted angle datasets, removing outliers and retaining 26 data points per target angle. At 120°, the system achieved an average accuracy of 5.26° with a precision of 3.26°, and at 20°, an accuracy of 7.11° with a precision of 4.01°.
These precision values were further interpreted in relation to a full 360° scale using the formula:
$\text{Precision}=\left(1-\frac{\text{Precision in Degrees}}{360}\right)\times 100$
This corresponds to approximately 98.9% and 98.8% accuracy, respectively, reinforcing the high reliability of our approach.
Table I: Accuracy and Precision at Different Angles
Metric 120 Degrees 20 Degrees
Number of Tests 30 30
Length of Trimmed Test 26 26
Accuracy (Degrees) 5.26 7.11
Precision (Degrees) 3.26 4.01
Precision Percentage 98.9% 98.8%
IV-D Conclusion
Overall, the results demonstrate that the proposed sound-tracking system is capable of estimating the direction of a sound source with high accuracy and consistency. The system remains robust under different angular positions and moderate environmental noise, making it suitable for applications such as smart home automation and security surveillance.
V Future Work
Based on the method presented in this study, there are several potential routes for future work to improve and expand the sound-tracking system’s capabilities. One possible enhancement would be to improve the precision of the sound-tracking by including more accurate microphones and implementing a beam-forming technique in addition to our approach to compute the direction of the sound source. Furthermore, incorporating machine learning techniques could improve the system’s ability to classify the direction of the sound based on audio signal characteristics, as well as anticipate the distance of the sound source.
Another direction for future work is to expand the capabilities of the system to track multiple sounds simultaneously by using multiple servo motors, each controlled by a separate set of microphones and vectors. Developing a user interface to display the direction of the sound and control the servo motor in real time could also be a valuable addition to the system.
To increase robustness, future versions of the system could integrate noise reduction methods and microphone calibration to mitigate environmental interference. The low-cost and low-complexity design has proven effective in controlled tests, and extending this to more complex environments—such as those with physical or acoustic obstacles—could help assess the system’s real-world performance and limitations.
Testing the system in various environments and scenarios may provide useful insights into its performance and highlight opportunities for improvement. Exploring the method’s possible applications in noise-canceling devices or security systems could also open up new avenues for this technology.
VI Conclusion
This paper presents a simple and cost-effective method for sound-tracking using an Arduino, three electret microphones, three LM358 op-amps, and a servo motor. By calculating the sum of vectors in rectangular form, we were able to determine the direction of the sound source with high accuracy. The results of our experiment demonstrated that the servo motor was able to track the direction of the sound source, and the recorded audio was clear and reliable.
The technique we presented in this study offers a practical and affordable solution for sound localization and tracking, with potential applications in various areas such as security, surveillance, and audio recording. It is particularly useful in scenarios where directional audio is required, such as in video conferencing, sound engineering, and audio equipment testing.
Direction detection plays a critical role in sound source localization, as it allows systems to spatially orient themselves toward a sound-emitting object. This directional awareness enhances the system’s ability to respond to audio stimuli, improves context awareness in robotics and automation, and enables more efficient resource focus in smart environments. Accurately detecting the direction from which a sound originates is often the first step in identifying, isolating, or responding to the source—making it a cornerstone for real-world applications.
Moreover, this method provides an excellent alternative to expensive commercial sound-tracking systems, which are often complex and difficult to use. It offers a simple and effective way to detect and record audio and control a servo motor based on the direction of the sound.
Finally, our study provides a valuable contribution to the field of sound-tracking and offers a promising solution for sound localization and tracking with an affordable and practical approach. Future research could focus on improving the design and implementation of the sound-tracking system by incorporating additional microphones and sensors to enhance its accuracy and precision. Furthermore, additional testing and experimentation could be conducted to evaluate the system’s performance in different environments and conditions.

### v3

@@
-Direction Estimation of Sound Sources Using Microphone Arrays and Signal Strength
+
 Abstract
@@
 Consequently, this technique presents a robust and reliable solution for sound-tracking and localization, with potential applications spanning diverse domains such as security systems, smart homes, and acoustic monitoring.
-Index Terms:
-Sound tracking, Sound source localization, Direction estimation, Microphone arrays
-The source code of the project is available on our GitHub repository:
-https://github.com/mahdi943/soundlocalization
-I Introduction
+The source code of the project is available on our GitHub repository: https://github.com/mahdi943/soundlocalization
+1 Introduction
 The ability to detect sounds and determine their direction has become important in modern intelligent systems.
@@
 Similarly, adaptive audio systems in home theaters use spatial awareness to enhance sound distribution, creating immersive listening experiences tailored to the user’s position.
-While sound localization generally refers to estimating the exact position of a sound source in space, sound tracking specifically focuses on determining the direction from which the sound originates.
-Localization aims to provide precise spatial coordinates, which often requires complex sensor arrays and extensive computation.
-In contrast, sound tracking emphasizes directional information, enabling real-time responsiveness with fewer sensors and simpler processing.
-This distinction allows sound tracking to serve as a lightweight, cost-effective alternative in scenarios where directionality is sufficient for the task at hand.
-Many existing sound localization methods rely on complex hardware setups, intensive signal processing, and high computational demands.
-While these approaches offer high accuracy, they are often impractical for lightweight, low-cost applications.
-To address this, we propose a simple and efficient method for detecting the direction of sound sources using a minimal sensor setup, designed to operate with limited processing resources.
-This work focuses on enabling directional sound-tracking for use in responsive systems such as smart home devices, interactive robots, and compact automation solutions.
-By estimating the direction of incoming sounds in real time, the proposed method supports adaptive behavior in resource-constrained environments, offering a practical alternative to more computationally intensive localization techniques.
-Our approach employs a vector-based technique to accurately detect sound direction by computing the sum of sound signal vectors arranged in a rectangular configuration.
-For instance, if a sound originates from the left, the vector representation will reflect this directional characteristic.
-This capability is crucial for swiftly identifying potential threats in security systems and enhancing the responsiveness of smart home devices by dynamically adjusting settings based on user positions.
-While machine learning and beamforming techniques are prevalent in the field of sound localization, our method distinguishes itself through its simplicity and affordability.
-By utilizing readily available Arduino boards and electret microphones, we simplify the assembly and configuration process, thereby enhancing accessibility for various applications.
-This approach holds great promise for improving systems that rely on precise sound source detection and tracking, especially in environments where budget constraints or deployment speed are critical considerations.
-II Related Work
-II-A Sound Source Localization (SSL)
+While sound localization estimates the exact position of a sound source, sound tracking focuses on determining its direction.
+Localization typically requires complex sensor arrays and heavy computation, whereas tracking provides directional information in real time with fewer sensors and simpler processing, offering a lightweight, cost-effective alternative when full coordinates are unnecessary.
+Many existing methods rely on complex hardware and intensive signal processing, which limits their practicality for low-cost applications.
+Our approach demonstrates that reliable direction estimation can be achieved using inexpensive, widely available hardware, without high-end processors or elaborate sensor arrays.
+This makes it suitable for educational, prototyping, and resource-constrained environments.
+This work focuses on enabling directional sound-tracking for responsive systems such as smart homes, robots, and compact automation tools.
+By estimating sound direction in real time, the method supports adaptive behavior in resource-constrained environments as a lightweight alternative to computationally intensive localization techniques.
+Our approach uses a vector-based method to compute the sum of microphone signal vectors, allowing accurate direction detection with minimal hardware.
+This capability is valuable for security, automation, and smart devices that benefit from rapid and reliable sound tracking.
+Unlike machine learning or beamforming approaches, our method emphasizes simplicity and affordability.
+Using readily available Arduino boards and electret microphones, we lower hardware costs and configuration effort, making sound-tracking accessible for budget-sensitive or rapid-deployment applications.
+2 Literture Review
+2.1 Sound Source Localization (SSL)
 Sound source localization (SSL) is a core component in many applications, including source separation [2], automatic speech recognition (ASR) [13], speech enhancement [27], noise control [3], and room acoustics analysis [1].
@@
 Recent research has explored machine learning approaches to classify sound direction using raw or pre-processed signal data [22, 19].
-II-B Sound Source Tracking
+In contrast to these hardware-heavy or computationally demanding approaches, our work emphasizes a lightweight solution that leverages low-cost, readily available components.
+This novelty allows us to achieve directional estimation without the complexity and resource requirements of traditional methods.
+2.2 Sound Source Tracking
 In dynamic environments, tracking the direction of a moving sound source over time is essential.
@@
 While less complex than TDOA or beamforming, these methods may offer faster processing at the expense of some spatial accuracy.
-II-C Summary and Motivation
+2.3 Summary and Motivation
 While traditional SSL and tracking methods offer high accuracy, they often require large microphone arrays, complex signal processing algorithms, and high computational power—making them less ideal for lightweight or low-cost systems.
-In contrast, our work focuses on a simple, intensity-based method that relies on only three microphones and does not require time delay or phase computations.
+Instead, our work focuses on a simple, intensity-based method that relies on only three microphones and does not require time delay or phase computations.
 This enables a more practical, scalable, and cost-effective solution for indoor directional sound-tracking, suitable for applications such as smart homes, speaker-following systems, and acoustic event detection.
-III Approach
-In this section, we describe the hardware and software setup used to implement our sound localization system.
-We then provide a detailed explanation of the components, the underlying methodology, and the measurement process involved in evaluating the system’s performance.
-III-A Methodology
+3 Methodology
 In this section, we explain how our approach determines the direction of a sound source by analyzing the intensity of sound signals captured by three microphones.
@@
 These samples reflect the sound signal’s strength over time and are used in the subsequent direction estimation process.
-1 Input: Sound signal $S$, number of samples NN, threshold $T$
+1Input: Sound signal $S$, number of samples NN, threshold $T$
 2 Output: Sound angle, sound magnitude, servo motor position
@@
 $\gamma=(240^{\circ},\text{PowerAvg}_{3})$ (3)
-These polar vectors are converted to rectangular (Cartesian) form using standard trigonometric conversion, resulting in vectors $R=(x,y)$, where $x=r\cdot\cos(\theta)$ and $y=r\cdot\sin(\theta)$.
+These polar vectors are converted by Equations 4, 5 and 6 to rectangular (Cartesian) form using standard trigonometric conversion, resulting in vectors $R=(x,y)$, where $x=r\cdot\cos(\theta)$ and $y=r\cdot\sin(\theta)$.
 $R_{1}=(r_{\alpha}\cdot\cos(\theta_{\alpha}),\ r_{\alpha}\cdot\sin(\theta_{\alpha}))$ (4)
@@
 $R_{3}=(r_{\gamma}\cdot\cos(\theta_{\gamma}),\ r_{\gamma}\cdot\sin(\theta_{\gamma}))$ (6)
-The three rectangular vectors are summed component-wise $R_{\text{total}}=(X,Y)$ to compute the resulting direction vector, where:
+The three rectangular vectors are summed component-wise $R_{\text{total}}=(X,Y)$ to compute the resulting direction vector, as defined in Equations 7 and 8.
 $X=\sum_{i=1}^{3}R_{x_{i}}$ (7)
 $Y=\sum_{i=1}^{3}R_{y_{i}}$ (8)
-The final direction of the sound source is determined by converting the resultant rectangular vector back to polar coordinates. The magnitude and angle are computed as follows:
+The final direction of the sound source is determined by converting the resultant rectangular vector back to polar coordinates. The magnitude and angle are computed as shown in Equations 9 and 10.
 $\text{ResultMag}=\sqrt{X^{2}+Y^{2}}$ (9)
@@
 Figure 1: Example showing the magnitudes of signals from three different microphones in black, and their vector sum in red, representing the resulting sound direction.
-III-B Implementation
+Figure 2: Circuit diagram represents the location of microphones and the servo motor in the center. Each microphone is 120 degrees from the next microphone, which shows on the green axis.
+3.1 Implementation
 The implemented sound-tracking system consists of three electret microphones arranged on a circular surface at equal angular intervals (0°, 120°, and 240°), each positioned 15 cm from the center.
 These microphones capture sound signals from different directions and are connected to LM358 operational amplifiers to boost the signal strength.
-The amplified outputs are fed into an Arduino board, which processes the signals and determines the direction of the incoming sound.
+The amplified outputs are fed into an Arduino Uno R3 microcontroller, which processes the signals and determines the direction of the incoming sound.
 A servo motor, also connected to the Arduino, rotates an indicator shaft toward the detected sound direction.
 Figure 2 illustrates the complete hardware setup, which is designed to be low-cost, efficient, and suitable for compact applications.
-Figure 2: Circuit diagram represents the location of microphones and the servo motor in the center. Each microphone is 120 degrees from the next microphone, which shows on the green axis.
-IV Results
-IV-A Experimental Setup and Design
+The low sampling rate (50 Hz) necessitated placing the sound source close to the microphone array (35 cm) to ensure reliable direction estimation.
+While this limits the testing range, it demonstrates the feasibility of ultra-low-cost hardware.
+4 Results
+4.1 Experimental Setup and Design
 To evaluate the performance of our sound-tracking system, we conducted 60 experiments using a 2-second audio clip played from a cellphone.
@@
 These were plotted as polar scatter plots (scatterpolar) to visually represent the direction and magnitude of the resulting vectors in (X,Y) format.
-IV-B Vector Visualization and Interpretation
+4.2 Vector Visualization and Interpretation
 For each angle setting (120° and 20°), 30 tests were conducted.
@@
 Figure 3: Detecting sound source angles and magnitude at 120 degrees
-Figure 4: Detecting sound source angles and magnitude at 0 degrees
+Figure 4: Detecting sound source angles and magnitude at 20 degrees
 Variability in signal magnitude arises due to environmental noise and the inherent sensitivity differences among microphones.
 These factors contribute to minor deviations from ideal alignment in the plotted vectors but do not significantly affect the overall directional accuracy.
-IV-C Accuracy and Precision Analysis
+4.3 Accuracy and Precision Analysis
 To ensure robustness, a 6% trimming was applied to the sorted angle datasets, removing outliers and retaining 26 data points per target angle.
@@
 This corresponds to approximately 98.9% and 98.8% accuracy, respectively, reinforcing the high reliability of our approach.
-Table I: Accuracy and Precision at Different Angles
+Table 1: Accuracy and Precision at Different Angles
 Metric 120 Degrees 20 Degrees
@@
 Precision Percentage 98.9% 98.8%
-IV-D Conclusion
-Overall, the results demonstrate that the proposed sound-tracking system is capable of estimating the direction of a sound source with high accuracy and consistency.
-The system remains robust under different angular positions and moderate environmental noise, making it suitable for applications such as smart home automation and security surveillance.
-V Future Work
-Based on the method presented in this study, there are several potential routes for future work to improve and expand the sound-tracking system’s capabilities.
-One possible enhancement would be to improve the precision of the sound-tracking by including more accurate microphones and implementing a beam-forming technique in addition to our approach to compute the direction of the sound source.
-Furthermore, incorporating machine learning techniques could improve the system’s ability to classify the direction of the sound based on audio signal characteristics, as well as anticipate the distance of the sound source.
-Another direction for future work is to expand the capabilities of the system to track multiple sounds simultaneously by using multiple servo motors, each controlled by a separate set of microphones and vectors.
-Developing a user interface to display the direction of the sound and control the servo motor in real time could also be a valuable addition to the system.
-To increase robustness, future versions of the system could integrate noise reduction methods and microphone calibration to mitigate environmental interference.
-The low-cost and low-complexity design has proven effective in controlled tests, and extending this to more complex environments—such as those with physical or acoustic obstacles—could help assess the system’s real-world performance and limitations.
-Testing the system in various environments and scenarios may provide useful insights into its performance and highlight opportunities for improvement.
-Exploring the method’s possible applications in noise-canceling devices or security systems could also open up new avenues for this technology.
-VI Conclusion
-This paper presents a simple and cost-effective method for sound-tracking using an Arduino, three electret microphones, three LM358 op-amps, and a servo motor.
-By calculating the sum of vectors in rectangular form, we were able to determine the direction of the sound source with high accuracy.
-The results of our experiment demonstrated that the servo motor was able to track the direction of the sound source, and the recorded audio was clear and reliable.
-The technique we presented in this study offers a practical and affordable solution for sound localization and tracking, with potential applications in various areas such as security, surveillance, and audio recording.
-It is particularly useful in scenarios where directional audio is required, such as in video conferencing, sound engineering, and audio equipment testing.
-Direction detection plays a critical role in sound source localization, as it allows systems to spatially orient themselves toward a sound-emitting object.
-This directional awareness enhances the system’s ability to respond to audio stimuli, improves context awareness in robotics and automation, and enables more efficient resource focus in smart environments.
-Accurately detecting the direction from which a sound originates is often the first step in identifying, isolating, or responding to the source—making it a cornerstone for real-world applications.
-Moreover, this method provides an excellent alternative to expensive commercial sound-tracking systems, which are often complex and difficult to use.
-It offers a simple and effective way to detect and record audio and control a servo motor based on the direction of the sound.
-Finally, our study provides a valuable contribution to the field of sound-tracking and offers a promising solution for sound localization and tracking with an affordable and practical approach.
-Future research could focus on improving the design and implementation of the sound-tracking system by incorporating additional microphones and sensors to enhance its accuracy and precision.
-Furthermore, additional testing and experimentation could be conducted to evaluate the system’s performance in different environments and conditions.
+5 Conclusion
+This paper introduced a simple and cost-effective method for sound-tracking using an Arduino, electret microphones, LM358 op-amps, and a servo motor.
+By applying a vector-sum approach, the system successfully estimated sound direction and controlled the servo to track the source.
+The results highlight the potential of this lightweight setup for practical applications in security, surveillance, robotics, and smart environments where directional audio is essential.
+The proposed method demonstrates an affordable alternative to complex and expensive sound-tracking systems, making it accessible for real-time use in constrained environments.
+While effective, the experiments were limited by the Arduino Uno’s low sampling rate and memory capacity, requiring the sound source to remain close to the microphone array.
+This design choice reflects our focus on validating the algorithm’s feasibility on low-resource hardware rather than optimizing for large-scale deployment.
+Future work will address these limitations by exploring higher sampling rates, more capable microcontrollers, and additional microphones or sensors to improve accuracy.
+Testing under realistic, noisy conditions will further evaluate robustness, extending the method’s applicability.
+Overall, this study contributes a practical foundation for affordable sound-tracking and opens opportunities for enhanced systems in robotics, automation, and acoustic sensing.